using the data in the table, determine the rate constant of the reaction and select the appropriate units.

A+2B -----> C+D

Table:

Trail / [A] (M) / [B] (M) / Rate (M/s)

1 / 0.220 / 0.370 / 0.0169

2 / 0.220 / 0.740 / 0.0169

3 / 0.440 / 0.370 / 0.0676


k=

Units=


Find k and units.



1
Expert's answer
2021-06-18T07:20:45-0400

Rate law V = K{A}a{B}b

V1/V2 = {A1/A2}a {B1/B2}b

Substituting trial 1 and trial 2 data in the equation above

0.0169/0.0169 = {0.22/0.22}a {0.74/0.37}b

1 = (1)a (2)b

20 = 2b since 1a = 1

b = 0

Similarly for trial 1 and 3

0.0676/0.0169 = (0.440/0.220)a (0.370/0.370)b

4 = (2)a (1)b

10 = (1)b since (2)a = 22

a = 2

V = K(A)2

Using trial data 1

0.0169 = K × (0.220)2

K = 0.0169/0.0484

K = 0.349M-1 S-1